OpenStudy (destinykiara99):

State the horizontal asymptote of the rational function. f(x) = quantity five x plus one divided by quantity nine x minus two.

jimthompson5910 (jim_thompson5910):

Just to be clear, the problem is this \[\Large f(x) = \frac{5x+1}{9x-2}\] right?

OpenStudy (destinykiara99):

Yes

OpenStudy (destinykiara99):

@jim_thompson5910 sorry, not sure why my page would not load

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

The degree of the numerator and the denominator are the same. They are both 1. Do you agree with this statement?

OpenStudy (destinykiara99):

Yes, I do

jimthompson5910 (jim_thompson5910):

since the degrees are the same, this means we just divide the leading coefficients which are 5 and 9 in this case, so 5/9

jimthompson5910 (jim_thompson5910):

the horizontal asymptote is simply the horizontal line y = 5/9 it's a horizontal line that cuts through 5/9 on the y axis

jimthompson5910 (jim_thompson5910):

5/9 = 0.5556 approximately

OpenStudy (destinykiara99):

Thank you so much!

jimthompson5910 (jim_thompson5910):

no problem